Mike has $1.55 in dimes and nickles. if he has 7 more nickels than dimes, how many of each coin does he have?

Answer by Paul(988)

Let the dimes be x
Let the nickels be x+7
Equation:
0.05(x+7)+0.1x=1.55
0.05x+0.35+0.1x=1.55
0.15x=1.2
x=8
8+7=15

Hence, their are 15 nickels and 15 dimes.
PAul.